4 pts How should the frequency of an oscillator be reported if its period is measured to be T 1.6 0.2 s? Note that the frequency is simply the inverse of the period, i.e., f-: 1 /T. Hint, the derivative of frequency with respect to the period, 7and the uncertainty of one quantity y that depends solely on one variable x is given by Ơy where the derivative is evaluated at the average value of a. Since frequency is the inverse of period, the units for frequency are inverse seconds, which are also called Hertz (abbreviate Hz). T2 0.63+0.39 Hz 0.63 ± 0.08 Hz 0.63 +0.2 Hz С 0.63 0.13 Hz
Answer #1

df/dT = -1/T2

for T = 1.6 s, df/dT = -1/1.62

so uncertainty in f is 0.2*1/1.62 = 0.08 Hz

hence second option is correct.